THE MILLIMETER- AND SUBMILLIMETER-WAVE SPECTRUM OF CYANOFORMAMIDE REVISITED

\underline{M. WINNEWISSER}, I. MEDVEDEV, F. C. DE LUCIA, E. HERBST, Department of Physics, The Ohio State University, Columbus, OH 43210; J. J. CHRISTIANSEN, Aakjars Alle 24B, 2860 Soborg, Denmark.